OpenStudy (anonymous):

simplify expression: -(-x+y-1)

OpenStudy (anonymous):

write it as -1(-x+y-1) Use Distributive Law, you get x - y + 1

jigglypuff314 (jigglypuff314):

distributive property can also be: a(b + c + d) = a*b + a*c + a*d so -1(-x + y -1) = (-1)(-x) + (-1)(y) + (-1)(-1) = x - y + 1 the concept is the same I like to think of it as what distributing means, like if I were to distribute bread I would break it into pieces and give it out so like if a was bread and b and c were who I gave it out to a(b + c), distributive property would be I give a bread to b and a bread to c so a*b + a*c

OpenStudy (anonymous):

ok...gotch...again much thanks

jigglypuff314 (jigglypuff314):

your welcome :)